# Bridging Online Behavior and Clinical Insight: A Longitudinal LLM-based Study of Suicidality on YouTube Reveals Novel Digital Markers

### v1

@@ 3.   Results
[…]
@@ 3.1 The bottom-up analysis
Our primary computational analysis examines LLM-generated topic distributions (Figure C.1). Statistical comparison between the treatment Attempted (During) group and the control groups found the two topics - YouTube Engagement and Mental Health Struggles, as both predictive of the pre-attempt period and significant indicators of temporal changes (pre to post-attempt). Representations of these topics are presented in Figure 7.
Figure 7: Word clouds generated from TF-IDF scores, highlighting the importance and frequency of words in the two topics that significantly predicted the pre-suicide attempt period and exhibited notable temporal changes across groups. The size of each word corresponds to its TF-IDF score. Below each topic, a representative transcript segment is presented.
@@ 3.1.1 What topics may serve as indicators of the period before a suicide attempt?
As part of the topic selection procedure, we identified topics predictive of the period preceding a suicide attempt with a forward-backward stepwise logistic regression [38], as these factors may act as indicators of current risk for suicide attempt.
The analysis revealed five significant topics (Table 3): (1) YouTube Engagement, (2) Mental Health Struggles, (3) Fashion and Style, (4) Rape and Sexual Assault and (5) LGBTQ+ Identity and Acceptance. Figure C.3 illustrates the word cloud of these topics, and Figure C.4 illustrates their temporal dynamics.
Treatment Group Control Group Topic $\beta$ Coeff. $z$ Score $p$ Value 95% CI Attempted (During) Attempted (Before) Mental Health Struggles -1.49 -3.62 $<0.001$ [-2.29, -0.68] Rape and Sexual Assault 0.68 2.68 0.007 [0.185, 1.186] Control (Major Life Event) YouTube Engagement -1.46 4.23 $<0.001$ [-2.142, -0.787] LGBTQ+ Identity and Acceptance 0.96 2.58 0.010 [0.233, 1.7] Control (Personal) YouTube Engagement -0.808 -3.19 0.001 [-1.304, -0.313] Fashion and Style -0.85 -2.74 0.006 [-1.462, -0.243] Combined Pool YouTube Engagement -0.73 -3.92 <0.001 [-1.098, -0.366]
Table 3: Regression Result: Each model predicts membership in the Attempted (During) group versus either an individual control group or a combination of all three (i.e., Combined Pool), using pre-event average topic probabilities as features. The analysis yielded four binary logistic regression models, each focusing on a different experimental setup.
@@ 3.1.2 Temporal effect analysis of pre-attempt indicator topics
Following the topic selection procedure (§3.1.1), which identified five topics, we conducted a two-way mixed ANOVA to examine the effects of group and time, as well as their interaction, across the four research groups. As shown in Table C.1 and illustrated in Figure 8, significant group-by-time interactions were observed for the topics YouTube Engagement and Mental Health Struggles. As illustrated in Figure 8, the Attempted (During) group exhibited a post-event increase in both topics, aligning with the Attempted (Before) group. For YouTube Engagement, both Attempted groups (Before and During) showed lower levels before and after the event compared to the control groups.
Figure 8: YouTube Engagement and Mental Health Struggles. The figures present the average topic values across four groups: Attempted (During), Attempted (Before), Control (Major Life Event), and Control (Personal), comparing Pre- and Post-Event (suicide attempt, major life event, or synthetic condition). The topics YouTube Engagement and Mental Health Struggles were first identified as significant in the bottom-up analysis (§3.1). The word clouds of these topics are presented in Figure 7.
These results were further examined using post-hoc Tukey’s HSD tests. Group-level differences were assessed via pairwise comparisons of overall topic means across groups (Table C.2), identifying specific group pairs with significant differences in topic prevalence. To investigate the group × time interaction effects, additional post-hoc analyses were conducted by comparing the temporal change (i.e., pre- vs. post-event differences) across groups (Table C.3). This analysis found that the Attempted (During) group showed significantly different temporal changes in YouTube Engagement compared to all control groups, and in Mental Health Struggles, only compared to the Attempted (Before) group.
@@ 3.2 The longitudinal topic analysis
Figure 9: Longitudinal Comparison of YouTube Engagement (left) and Mental Health Struggles (right) relative to $t=0$ in the Attempted (During) Group. The figures depict temporal trends of these topics within the Attempted (During) group, with the timeline aligned to the suicide attempt. At each time step, comparisons were made to $t=0$ using the paired t-test. Asterisks (*) indicate statistically significant differences. The full statistical procedure is detailed in §B.3.
@@
Figure 9 shows the results of the within-group analysis, which illustrates how the prevalence of YouTube Engagement and Mental Health Struggles significantly increased after the attempt. Additionally, the between-group analysis, depicted in Figure 10, reveals that YouTube Engagement significantly differs between the Attempted (During) group and all control groups in the time-steps right before the attempt. Meanwhile, significant differences in Mental Health Struggles were primarily observed between the Attempted (During) and Attempted (Before) groups. These differences were further validated as relative to suicide attempt, as a control analysis using COVID-19 as a reference event showed no significant differences between-group effects (§C.3.3).
Figure 10: Longitudinal Comparison of YouTube Engagement (left) and Mental Health Struggles (right) Between Attempted (During) and Control Groups. The figures illustrate temporal trends of these topics across the four groups. The timeline is aligned to a unified event-based reference point (suicide attempt, major life event, or synthetic condition). At each time step, the Attempted (During) group was compared to the other groups using the t-test. Asterisks (*) denote statistically significant differences. The full statistical procedure is detailed in §B.3
@@ 3.3 The top-down domain-expert analysis
We found that the Attempted (During) and Attempted (Before) groups reported similar risk and resilience factors in their suicide narratives (Table 4). Correspondingly, the users in both the Attempted (During) and the Attempted (Before), display a similar distribution of mental health conditions (e.g., 78.6% and 79.9% self-reported depression), resilience factors (40.5% and 37.4% self-reported social/family support), and experiencing feelings of hopelessness/despair (71.4% and 64%) and unbearable mental pain (38.1% and 46.8%). The only significant difference between the groups lies in their attempt sharing motivations: Attempted (During) users mainly discuss Personal recovery (66.7% vs. 38.1%), while Attempted (Before) users mentioned the Desire to help others (77% vs. 40.3%).
[…]

### v2

@@ 3.  Results
Table 1: Comparison of analytical approaches and suicide-related factors. Each factor was identified through one or more of the following approaches: Bottom-Up $\uparrow$ (LM-based topic modeling), Hybrid $\uparrow\downarrow$ (expert evaluation of LM-based topics), or Top-Down $\downarrow$ (psychological annotation of narrative videos). The leftmost column indicates whether the factor was generated by an LLM or defined by a human expert. The rightmost columns indicate statistically significant effects from the mixed ANOVA: Group Effect refers to the main effect of group; Temporal Effect refers to the group × time interaction. "–" marks untested effects. Attempted (During) group.
Origin Factor Bottom-Up $\uparrow$ Hybrid $\uparrow\downarrow$ Top-Down $\downarrow$ Group Temporal
LM-Based Topics YouTube Engagement
Mental Health Struggles
Fashion and Style
LGBTQ+ and Personal Identity
Rape and Sexual Assault
Domain-Expert Personal Recovery Process
Desire to Help Others
@@ 3.1 The bottom-up analysis
Our primary computational analysis examines LM-based topic distributions (Figure B.1). Statistical comparison between the treatment Attempted (During) group and the control groups found two topics - YouTube Engagement and Mental Health Struggles, as both predictive of the pre-attempt period and significant indicators of temporal changes (pre to post-attempt). Representations of these topics are shown in Figure 4.
Figure 4: YouTube Engagement and Mental Health Struggles: The figures present topics identified as significant in the bottom-up analysis (§3.1): YouTube Engagement (left) and Mental Health Struggles (right). The top row shows average topic values across four groups—Attempted (During), Attempted (Before), Control (Major Life Event), and Control (Personal)—in Pre- and Post-Event periods. TF-IDF word clouds are displayed alongside representative transcript excerpts, with word size reflecting importance. Below, the second row presents longitudinal trends over normalized time, aligned to each group’s reference event (e.g., suicide attempt). Asterisks (*) denote statistically significant differences between Attempted (During) and other groups at each time point (t-test; see Appendix B.3 for temporal testing procedures).
@@ 3.1.1 What topics may serve as indicators of the period before a suicide attempt?
As part of the topic selection procedure, we identified topics predictive of the period preceding a suicide attempt with a forward-backward stepwise logistic regression4, as these factors may act as indicators of current risk for suicide attempt.
The analysis revealed five significant topics (Table 2): (1) YouTube Engagement, (2) Mental Health Struggles, (3) Fashion and Style, (4) Rape and Sexual Assault and (5) LGBTQ+ Identity and Acceptance. Figures C.3 and C.4 illustrate their word cloud and temporal dynamics.
Table 2: Regression results predicting membership in the Attempted (During) group versus individual or pooled control groups (Four separate models were trained). Models use pre-event average topic probabilities. $R^{2}$ refers to McFadden’s pseudo-R-squared
Treatment Group Control Group Topic $\beta$ Coeff. $z$ Score $p$ Value 95% CI $R^{2}$
Attempted (During) Attempted (Before) Mental Health Struggles -1.49 -3.62 <0.001<0.001 [-2.29, -0.68] 0.11
Rape and Sexual Assault 0.68 2.68 0.007 [0.185, 1.186]
Control (Major Life Event) YouTube Engagement -1.46 4.23 <0.001<0.001 [-2.142, -0.787] 0.23
LGBTQ+ Identity and Acceptance 0.96 2.58 0.010 [0.233, 1.7]
Control (Matches) YouTube Engagement -0.808 -3.19 0.001 [-1.304, -0.313] 0.22
Fashion and Style -0.85 -2.74 0.006 [-1.462, -0.243]
Combined Pool YouTube Engagement -0.73 -3.92 <0.001<0.001 [-1.098, -0.366] 0.05
@@ 3.1.2 Temporal effect analysis of pre-attempt indicator topics
Table 3: Fixed effects from GLMMs for topic‐level outcomes. Positive coefficients indicate higher topic prevalence relative to baseline. Reference group: Attempted (During); Reference time: Before event.
Topic Parameter $\beta$ OR SE z p $p_{adj}$
Temporal Effects (Time and Group × Time Interactions)
Mental Health Struggles Time (After vs. Before) 0.554 1.74 0.145 3.82 <.001<.001 <.001<.001 **
Group × Time: Attempted (Before) 0.440-0.440 0.64 0.162 2.72-2.72 .006 .016 *
Group × Time: Control (Matches) 0.356-0.356 0.70 0.172 2.08-2.08 .038 .095
Group × Time: Control (Major Life Event) 0.397-0.397 0.67 0.199 1.99-1.99 .046 .11
YouTube Engagement Time (After vs. Before) 0.515 1.67 0.107 4.80 <.001<.001 <.001<.001 **
Group × Time: Attempted (Before) 0.353-0.353 0.70 0.122 2.90-2.90 .004 .016 *
Group × Time: Control (Matches) 0.527-0.527 0.59 0.128 4.13-4.13 <.001<.001 <.001<.001 **
Group × Time: Control (Major Life Event) 0.415-0.415 0.66 0.149 2.78-2.78 .005 .027 **
Non‐Temporal Effects (Baseline and Covariates)
Fashion and Style Group: Control (Matches) 0.891 2.44 0.311 2.86 .004 .01 **
Gender (Male) 0.661-0.661 0.52 0.259 2.55-2.55 .011 .026 *
YouTube Engagement Group: Attempted (Before) 0.336 1.40 0.138 2.43 .015 .075
Group: Control (Matches) 0.827 2.29 0.137 6.05 <.001<.001 <.001<.001 **
Group: Control (Major Life Event) 0.764 2.15 0.204 3.75 <.001<.001 .001 **
Gender (Male) 0.314-0.314 0.73 0.110 2.84-2.84 .005 .023 *
Coefficients ($\beta$) are log‐odds from Beta GLMMs; odds ratios (OR) are computed as $\exp(\beta)$. $p$-values are reported for each model coefficient. To account for multiple comparisons across topics, $p_{adj}$ shows the FDR-corrected p-values (Benjamini-Hochberg) applied separately for each coefficient type across all topics. Significance stars correspond to $p_{adj}$ thresholds: * <0.05<0.05, ** <0.01<0.01.
Following the topic selection procedure (§3.1.1), which identified five topics, we fit a generalized linear mixed model (GLMM) to examine the effects of group and time, as well as their interaction, along with demographic and upload activity covariates.
As illustrated in Figure 4, the Attempted (During) group exhibited a post-event increase in both topics, aligning with the Attempted (Before) group. For YouTube Engagement, both Attempted groups (Before and During) showed lower levels before and after the event compared to the control groups.
As shown in Table 3 and illustrated in Figure 4, significant group-by-time interactions were observed for the topics YouTube Engagement and Mental Health Struggles. This analysis found that the Attempted (During) group showed significantly different temporal changes in YouTube Engagement compared to all control groups, and in Mental Health Struggles, only compared to the Attempted (Before) group. We provide a more detailed description of the GLMM model and results, in terms of odds change and non-temporal significant results in Appendix D.1 (Effect Analysis). We also conducted a sensitivity analysis excluding ambiguous cases (Appendix D.2), which confirmed that temporal effects remained statistically significant and consistent in direction.
@@ 3.2 The longitudinal topic analysis
Figure C.6 shows the results of the within-group analysis, which illustrates how the prevalence of YouTube Engagement and Mental Health Struggles significantly increased after the attempt. Additionally, the between-group analysis (Figure 4) reveals that YouTube Engagement significantly differs between the Attempted (During) group and all control groups in the time-steps right before the attempt. Meanwhile, significant differences in Mental Health Struggles were primarily observed between the Attempted (During) and Attempted (Before) groups. These differences were further validated as relative to suicide attempt, as a control analysis using COVID-19 as a reference event showed no significant differences between-group effects (Appendix D.2).
@@ 3.3 The top-down domain-expert analysis
We found that the Attempted (During) and Attempted (Before) groups reported similar risk and resilience factors in their suicide narratives (Table C.1). Correspondingly, the users in both the Attempted (During) and the Attempted (Before), display a similar distribution of mental health conditions (e.g., 78.6% and 79.9% self-reported depression), resilience factors (40.5% and 37.4% self-reported social/family support), and feelings of hopelessness/despair (71.4% and 64%) and unbearable mental pain (38.1% and 46.8%). The only significant difference between the groups lies in their attempt sharing motivations: Attempted (During) users mainly discuss Personal recovery (66.7% vs. 38.1%), while Attempted (Before) users mentioned the Desire to help others (77% vs. 40.3%).
Subsequently, we examined the significance of these differences by applying one-hot encoding to the psychological factors and conducting a stepwise logistic regression analysis on group membership. The regression identified the motivation Desire to help others, as a significant factor ($\beta=-1.69$, $p<.01$, z=4.49z=-4.49), followed by Personal recovery process ($\beta=1.08$, $p<.01$, $z=2.93$) after the former was excluded. These were the only two significant factors; excluding both revealed no additional factors (see §4).